OpenStudy (anonymous):

Jamie is 5 years older than Ella. Jamie's age is 11 years less than three times Ella's age. The system below models the relationship between Jamie's age (j) and Ella's age (e): j = e + 5 j = 3e − 11 Which of the following methods is correct to find Jamie's and Ella's age? Solve j + 5 = 3j − 11 to find the value of j. Write the points where the graphs of the equations intersect the x-axis. Write the points where the graphs of the equations intersect the y-axis. Solve e + 5 = 3e − 11 to find the value of e. I chose D, am I right ?

OpenStudy (triciaal):

there are different ways of expressing the same value example 10 = 4 + 6 and 10 = 8 + 2 and 10 = 7 + 3 4 + 6 = 8 + 2 = 7 + 3 all = 10 using this concept same applies to variables j = j so e + 5 = 3e - 1 there is one equation with one unknown variable

OpenStudy (triciaal):

yes D is correct

OpenStudy (anonymous):

Ok, thank you

OpenStudy (triciaal):

welcome